# Positive concave deep equilibrium models

### v1

@@ Abstract
Deep equilibrium (DEQ) models are widely recognized as a memory efficient alternative to standard neural networks, achieving state-of-the-art performance in language modeling and computer vision tasks. These models solve a fixed point equation instead of explicitly computing the output, which sets them apart from standard neural networks. However, existing DEQ models often lack formal guarantees of the existence and uniqueness of the fixed point, and the convergence of the numerical scheme used for computing the fixed point is not formally established. As a result, DEQ models are potentially unstable in practice. To address these drawbacks, we introduce a novel class of DEQ models called positive concave deep equilibrium (pcDEQ) models. Our approach, which is based on nonlinear Perron-Frobenius theory, enforces nonnegative weights and activation functions that are concave on the positive orthant. By imposing these constraints, we can easily ensure the existence and uniqueness of the fixed point without relying on additional complex assumptions commonly found in the DEQ literature, such as those based on monotone operator theory in convex analysis. Furthermore, the fixed point can be computed with the standard fixed point algorithm, and we provide theoretical guarantees of geometric convergence, which, in particular, simplifies the training process. Experiments demonstrate the competitiveness of our pcDEQ models against other implicit models.
@@ 1 Introduction
Implicit models (Bai et al., 2019, 2020; Chen et al., 2018; El Ghaoui et al., 2021; Baker et al., 2023; Tsuchida & Ong, 2023) are attracting considerable interest owing to their improved memory efficiency compared to standard neural networks. These models solve implicit equations instead of explicitly computing the output of the layers, and they can be divided into two main categories: neural ordinary differential equations (Neural ODEs) (Chen et al., 2018) and deep equilibrium models (DEQ) models (Bai et al., 2019).
Neural ODEs solve differential equations parameterized by the neural network input, with the output representing the solution to these equations. On the other hand, the implicit layers of DEQ models solve nonlinear fixed point equations that are not necessarily derived from differential equations. An interesting aspect about DEQ models is that a single implicit DEQ layer emulates a standard neural network with an infinite number of layers and tied weights. While both DEQ models and neural ODEs require constant memory training, DEQ models often outperform neural ODEs, achieving state-of-the-art results in language modeling tasks (Bai et al., 2019) and computer vision tasks (Bai et al., 2020; Xie et al., 2022). However, a potential limitation of standard DEQ models is that they are based on iterative methods that require careful initialization, tuning of hyperparameters, and special regularization (e.g., recurrent dropout) to ensure convergence to the fixed point, which is hard to guarantee with existing approaches. In general, standard DEQ models operate heuristically, and in many cases they even lack formal guarantees regarding the existence and uniqueness of the fixed point problem being solved.
@@
    We propose a new class of DEQ models, called pcDEQ models, which are based on nonlinear Perron-Frobenius theory and are equipped with guarantees of the existence and uniqueness of the fixed point of pcDEQ models. Furthermore, we prove that the standard fixed point iteration for the proposed pcDEQ models converges to the fixed point geometrically fast.
@@
    We show that pcDEQ models are based on easily verifiable assumptions needed to establish the existence and uniqueness of the fixed point.
  4. 4.
    We demonstrate the competitiveness of the proposed approach in terms of accuracy and number of parameters over existing alternatives for image classification tasks.
@@ 2 Related work
In the seminal paper (Bai et al., 2019), DEQ models have been applied to language modeling tasks, and they have been shown to outperform standard neural networks constructed with a similar number of parameters. In subsequent studies, Bai et al. proposed multiscale extensions of DEQ models (Bai et al., 2020), where a single DEQ model is used for image classification and image segmentation. Since these pioneering studies, DEQ models have been successful in many applications, including, to name a few, object detection (Wang et al., 2023), optical flow estimation (Bai et al., 2022), video semantic segmentation (Ertenli et al., 2022), medical image segmentation (Zhang et al., 2022), snapshot compressive imaging (Zhao et al., 2023), and image denoising (Chen et al., 2023; Gkillas et al., 2023a), machine translation (Zheng et al., 2023), inverse problems (Gilton et al., 2021; Zou et al., 2023), music source separation (Koyama et al., 2022), federated learning (Gkillas et al., 2023b), and diffusion models (Geng et al., 2023; Pokle et al., 2022). These models are strongly based on the fixed point theory that we summarize below.
Let $\mathbb{R}^{n}$ be the Euclidean space and let $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$. If $f$ is a Lipschitz contraction, meaning it is Lipschitz continuous with a Lipschitz constant $L<1$, then the Banach fixed point theorem guarantees the uniqueness and existence of the fixed point $x^{\star}=f(x^{\star})$ in $\mathbb{R}^{n}$. Moreover, the fixed point iteration $x_{k+1}=f(x_{k})$ converges linearly to $x^{\star}$ for any initial point $x_{1}\in\mathbb{R}^{n}$. However, if $f$ is not a Lipschitz contraction, questions related to the existence and uniqueness of the unique point, and also the convergence of the fixed point iteration, become more delicate, and these issues have been the subject of extensive research in the mathematical literature.
In modern convex analysis, there has been a significant focus on nonexpansive mappings in Hilbert spaces (Yamada et al., 2011), which are mappings with Lipschitz constant equal to one, and their relation to monotone operator theory (Bauschke & Combettes, 2017). Nonexpansive mappings do not necessarily have a fixed point, and, if the fixed point set is nonempty, it may not be a singleton in general. Furthermore, if the fixed point exists, then the fixed point iteration may fail to converge, in which case we can use the Krasnosel’skii-Mann iteration. Particular instances of this iteration include well-known algorithms used in machine learning, such as the projected gradient method, the proximal forward-backward splitting method, the Douglas-Rachford splitting method, the projection onto convex sets method, and many others (Yamada et al., 2011).
@@
To avoid the above difficulties, we use in this study nonlinear Perron-Frobenius theory (Krause, 2015; Lemmens & Nussbaum, 2012) as an alternative to traditional theory in Hilbert spaces. More specifically, we consider deep equilibrium layers that belong to the class of standard interference mappings. These mappings, introduced in the next section, have been widely used in wireless networks (Yates, 1995; Schubert & Boche, 2011; Stanczak et al., 2009; You & Yuan, 2020; Miretti et al., 2023a, b; Shindoh, 2020, 2019; Cavalcante et al., 2015, 2019). They are not necessarily nonexpansive in Hilbert spaces, but they are contractive (though not necessarily Lipschitz contractions) in metric spaces defined on cones, such as the cone of positive vectors. Before introducing SI mappings and their applications in DEQ models, we first need to establish the notation and formally introduce the concept of DEQ layers.
@@ 3 Preliminaries
Let $x,y\in\mathbb{R}^{n}_{+}$. The partial ordering induced by the nonnegative cone is denoted as $x\leq y\Leftrightarrow y-x\in\mathbb{R}^{n}_{+}$. In a similar way, for $x\neq y$, $x<y\Leftrightarrow y-x\in\mathbb{R}^{n}_{+}$, and $x\ll y\Leftrightarrow y-x\in\operatorname{int}(\mathbb{R}^{n}_{+})$. The fixed point set of a function $f:X\rightarrow Y$ with $Y\subset X$ is denoted as
@@ 3.1 Deep equilibrium layers
We now have all the necessary notation to introduce DEQ models.
@@
A DEQ layer consists of an implicit function mapping an input $x\in X\subset\mathbb{R}^{n}$ to an output $z^{\star}\in\textup{Fix}(g_{x})\subset Y\subset X$, where $g_{x}:X\to Y$ is an explicit function given by
@@
$W\colon X\to Y$ is a linear operator (weights); and $\sigma\colon X\to Y$ is a (vector-valued nonlinear) activation function, composed elementwise from a given scalar activation function.
In the above definition, closed-form expressions for the implicit function $x\mapsto z^{\star}\in\textup{Fix}(g_{x})$ are not required, and the numerical scheme used to compute the output $z^{\star}\in\textup{Fix}(g_{x})$ from a given input $x$ is not specified. As a result, to ensure that the implicit function $x\mapsto z^{\star}\in\textup{Fix}(g_{x})$ is well-defined and independent of the numerical scheme used to compute the output, we require $\textup{Fix}(g_{x})$ to be a singleton for every $x\subset X$. This fixed point formulation allows for direct implicit differentiation, which is crucial for training DEQ models (Bai et al., 2019). In the text that follows, for convenience, we always refer to a DEQ layer using its explicit function $g_{x}$ because, with the above restriction of $\mathrm{Fix}(g_{x})$ being a singleton, the implicit function $x\mapsto z^{\star}\in\mathrm{Fix}(g_{x})$ is uniquely determined.
One of the simplest numerical schemes for computing fixed points, which is the numerical scheme we consider in this study to avoid unnecessary technical digressions, is the standard fixed point iteration, which, using the notation in Definition 3.1, we can write as
@@ 3.2 Standard interference and positive concave mappings
    $(\forall x\in\mathbb{R}^{n}_{+})(\forall\lambda>1)\quad g(\lambda x)\ll g(x).$ (4)
@@
(We emphasize that NC mappings are not SI mappings in general, only PC mappings are guaranteed to have this property without any further assumptions.)
@@ 4 PC deep equilibrium layers
We recall that the activation functions from the above lists are applied elementwise to vector arguments, see Definition 3.1. The following lemma provides two simple and sufficient conditions to construct positive concave DEQ layers.
@@
Consider a DEQ layer $g_{x}\colon X\to Y$ of the form in (1) for an input $x$. Then:
  • (Assumption 1) $z\mapsto g_{x}(z):=\sigma(Wz+x)$ is a PC mapping $g_{x}\colon\mathbb{R}^{n}_{+}\to\operatorname{int}(\mathbb{R}^{n}_{+})$ if $W\in\mathbb{R}^{n\times n}_{+}$, $x\in\textup{int}(\mathbb{R}^{n}_{+}),z\in\mathbb{R}^{n}_{+}$, and $\sigma$ is constructed elementwise from any scalar activation function from List 1 in Remark 4.1;
  • (Assumption 2) $z\mapsto g_{x}(z):=\sigma(Wz+x)$ is a PC mapping $g_{x}\colon\mathbb{R}^{n}_{+}\to\operatorname{int}(\mathbb{R}^{n}_{+})$ if $W\in\mathbb{R}^{n\times n}_{+}$, $x\in\mathbb{R}^{n}_{+},z\in\mathbb{R}^{n}_{+}$, and $\sigma$ is constructed elementwise from the scalar activation function from List 2 in Remark 4.1.
@@
On the other hand, if Assumption 2 is satisfied, we note that, in this case, $\sigma(0)\gg 0$, hence $g_{x}(z)=(\sigma\circ f_{x})(z)$ for all $z\in\mathbb{R}^{n}_{+}$, which implies $g_{x}$ is also a PC mapping in this case. \ $\blacksquare$
@@
Let $g_{x}\colon\mathbb{R}^{n}_{+}\mapsto\textup{int}(\mathbb{R}^{n}_{+})$ be a DEQ layer of the form in (1) for a given input $x$. If $g_{x}$ (including the input $x$) satisfies Assumption 1 in Lemma 4.2, then $g_{x}$ is called a pcDEQ-1 layer. If $g_{x}$ satisfies Assumption 2 in Lemma 4.2, then $g_{x}$ is called a pcDEQ-2 layer. By pcDEQ we mean a layer $g_{x}$ that is either a pcDEQ-1 layer or a pcDEQ-2 layer.
@@
Figure 1: The visualization of the possible construction of pcDEQ layers. The symbols shown in the figures mean: $W_{+}$ nonnegative weights, $z_{+}$ nonnegative vector of fixed point iteration, $x_{++}$ positive input to the layer, $\sigma_{NC}$ nonnegative concave activation function (List 1 in Remark 4.1) and $\sigma_{PC}$ positive concave activation function (List 2 in Remark 4.1).
Lemma 4.2 asserts that pcDEQ layers are PC mappings for any allowed input $x$, which, in view of Proposition A.6, implies that they are also SI mappings. We can now use Definition A.2 to provide the form of the asymptotic mapping associated with a pcDEQ layer. The assertion of Proposition 4.4 follows from Proposition 11 in (Piotrowski et al., 2021), and, for completeness, we include the proof below.
@@
$g_{x,\infty}(z_{k})=\lim_{p\to\infty}\sigma^{\prime}(Wpz+x)Wz.$ (7)
For activation functions composed from any scalar activation functions in Remark 4.1, we have $\lim_{p\to\infty}\sigma^{\prime}(Wpz)=0$, hence also
@@
Choose any input $x$ satisfying the conditions in Definition 4.3. From Lemma 4.2, $g_{x}$ is a PC mapping. Thus, from Proposition A.6, $g_{x}$ is also an SI mapping. From Corollary 4.5, the nonlinear spectral radius of $g_{x}$ is $p(g_{x,\infty})=0.$ If follows from Proposition A.4 that $g_{x}$ has a unique fixed point $z^{\star}.$ Then, from Proposition A.7 it follows that the fixed point iteration of $g_{x}$ in (2) converges geometrically to $z^{\star}$ with a factor $c\in[0,1)$ for any starting point $z_{1}\in\mathbb{R}^{n}_{+}.$blacksquare

### v2

@@ Abstract
Deep equilibrium (DEQ) models are widely recognized as a memory efficient alternative to standard neural networks, achieving state-of-the-art performance in language modeling and computer vision tasks. These models solve a fixed point equation instead of explicitly computing the output, which sets them apart from standard neural networks. However, existing DEQ models often lack formal guarantees of the existence and uniqueness of the fixed point, and the convergence of the numerical scheme used for computing the fixed point is not formally established. As a result, DEQ models are potentially unstable in practice. To address these drawbacks, we introduce a novel class of DEQ models called positive concave deep equilibrium (pcDEQ) models. Our approach, which is based on nonlinear Perron-Frobenius theory, enforces nonnegative weights and activation functions that are concave on the positive orthant. By imposing these constraints, we can easily ensure the existence and uniqueness of the fixed point without relying on additional complex assumptions commonly found in the DEQ literature, such as those based on monotone operator theory in convex analysis. Furthermore, the fixed point can be computed with the standard fixed point algorithm, and we provide theoretical guarantees of its geometric convergence, which, in particular, simplifies the training process. Experiments demonstrate the competitiveness of our pcDEQ models against other implicit models.
@@ 1 Introduction
Implicit models (Bai et al., 2019, 2020; Chen et al., 2018; El Ghaoui et al., 2021; Baker et al., 2023; Tsuchida & Ong, 2023; Revay et al., 2020; Wei & Kolter, 2021; Geng et al., 2021) are attracting considerable interest owing to their improved memory efficiency compared to standard neural networks. These models solve implicit equations instead of explicitly computing the output of the layers, and they can be divided into two main categories: neural ordinary differential equations (Neural ODEs) (Chen et al., 2018) and deep equilibrium models (DEQ) models (Bai et al., 2019).
Neural ODEs solve differential equations parameterized by the neural network input, with the output representing the solution to these equations. On the other hand, the implicit layers of DEQ models solve nonlinear fixed point equations that are not necessarily derived from differential equations. An interesting aspect about DEQ models is that a single implicit DEQ layer emulates a standard neural network with an infinite number of layers and tied weights. While both DEQ models and neural ODEs require constant training memory, DEQ models often outperform neural ODEs, achieving state-of-the-art results in language modeling tasks (Bai et al., 2019) and computer vision tasks (Bai et al., 2020; Xie et al., 2022). However, a potential limitation of standard DEQ models is that they are based on iterative methods that require careful initialization, tuning of hyperparameters, and special regularization (e.g., recurrent dropout) to ensure convergence to the fixed point, which is hard to guarantee with existing approaches. In general, standard DEQ models operate heuristically, and in many cases they even lack formal guarantees regarding the existence and uniqueness of the fixed point problem being solved.
@@
    We propose a new class of DEQ models, called pcDEQ models, which are based on nonlinear Perron-Frobenius theory and are equipped with guarantees of the existence and uniqueness of the fixed point. Furthermore, we prove that the standard fixed point iteration for the proposed pcDEQ models converges to the fixed point geometrically fast.
@@
    We demonstrate competitive improvement of the proposed approach in terms of accuracy and number of parameters over existing alternatives for image classification tasks.
@@ 2 Related Work
In the seminal paper (Bai et al., 2019), DEQ models have been applied to language modeling tasks, and they have been shown to outperform standard neural networks constructed with a similar number of parameters. In subsequent studies, Bai et al. proposed multiscale extensions of DEQ models (Bai et al., 2020), where a single DEQ model is used for image classification and image segmentation. Since these pioneering studies, DEQ models have been successful in many applications, including, to name a few, object detection (Wang et al., 2023), optical flow estimation (Bai et al., 2022), video semantic segmentation (Ertenli et al., 2022), medical image segmentation (Zhang et al., 2022), snapshot compressive imaging (Zhao et al., 2023), image denoising (Chen et al., 2023; Gkillas et al., 2023a), machine translation (Zheng et al., 2023), inverse problems (Gilton et al., 2021; Zou et al., 2023), music source separation (Koyama et al., 2022), federated learning (Gkillas et al., 2023b), and diffusion models (Geng et al., 2023; Pokle et al., 2022). These models are strongly based on the fixed point theory that we summarize below.
Let $\mathbb{R}^{n}$ be the standard Euclidean metric space and let $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$. If $f$ is a Lipschitz contraction, meaning it is Lipschitz continuous with a Lipschitz constant $L<1$, then the Banach fixed point theorem guarantees the uniqueness and existence of the fixed point $x^{\star}=f(x^{\star})$ in $\mathbb{R}^{n}$. Moreover, the fixed point iteration $x_{k+1}=f(x_{k})$ converges linearly to $x^{\star}$ for any initial point $x_{1}\in\mathbb{R}^{n}$. However, if $f$ is not a Lipschitz contraction, questions related to the existence and uniqueness of the unique point, and also the convergence of the fixed point iteration, become more delicate, and these issues have been the subject of extensive research in the mathematical literature.
In modern convex analysis, there has been a significant focus on nonexpansive mappings in Hilbert spaces (Yamada et al., 2011), which are mappings with Lipschitz constant equal to one, and their relation to monotone operator theory (Bauschke & Combettes, 2017). Nonexpansive mappings do not necessarily have a fixed point, and, if the fixed point set is nonempty, it may not be a singleton in general. Furthermore, even if the fixed point exists, then the fixed point iteration may fail to converge, in which case one can resort to various iterative methods based on the the Krasnosel’skii-Mann iteration to ensure it. Particular instances of this iteration include well-known algorithms used in machine learning, such as the projected gradient method, the proximal forward-backward splitting method, the Douglas-Rachford splitting method, the projection onto convex sets method, and many others (Yamada et al., 2011).
@@
To avoid the above difficulties, we use in this study nonlinear Perron-Frobenius theory (Krause, 2015; Lemmens & Nussbaum, 2012) as an alternative to traditional theory in Hilbert spaces. More specifically, we consider deep equilibrium layers that belong to the class of standard interference mappings. These mappings, introduced in the next section, have been widely used in wireless networks (Yates, 1995; Schubert & Boche, 2011; Stanczak et al., 2009; You & Yuan, 2020; Miretti et al., 2023a, b; Shindoh, 2020, 2019; Cavalcante et al., 2016, 2019). They are not necessarily nonexpansive in Hilbert spaces, but they are contractive (though not necessarily Lipschitz contractions) in some metric spaces defined on cones, such as the cone of positive vectors. Before introducing SI mappings and their applications in DEQ models, we first need to establish the notation and formally introduce the concept of DEQ layers.
@@ 3 Preliminaries
Let $x,y\in\mathbb{R}^{n}_{+}$. The partial ordering induced by the nonnegative cone is denoted as $x\leq y\Leftrightarrow y-x\in\mathbb{R}^{n}_{+}$. In a similar way, for $x\neq y$, $x<y\Leftrightarrow y-x\in\mathbb{R}^{n}_{+}$, and $x\ll y\Leftrightarrow y-x\in\operatorname{int}(\mathbb{R}^{n}_{+})$. The fixed point set of a function $f:X\rightarrow Y$ with $Y$ and $X$ being subsets of a given set $S$ is denoted as
@@ 3.1 Deep Equilibrium Layers
We now have all the necessary notation to introduce generic DEQ models.
@@
A DEQ layer maps an input $x\in X\subset\mathbb{R}^{n}$ to an output $z^{\star}\in\textup{Fix}(g_{x})\subset Y\subset X$, where $g_{x}:X\to Y$ is an explicit function given by
@@
$W\colon X\to Y$ is a linear operator (weight matrix); and $\sigma\colon X\to Y$ is a (vector-valued nonlinear) activation function, composed elementwise from a given scalar activation function.
In the above definition, closed-form expressions for the implicit function $x\mapsto z^{\star}\in\textup{Fix}(g_{x})$ are not required, and the numerical scheme used to compute the output $z^{\star}\in\textup{Fix}(g_{x})$ from a given input $x$ is not specified. As a result, to ensure that the implicit function $x\mapsto z^{\star}\in\textup{Fix}(g_{x})$ is well-defined and independent of the numerical scheme used to compute the output, we require $\textup{Fix}(g_{x})$ to be a singleton for every $x\subset X$. This fixed point formulation allows for direct implicit differentiation, which is crucial for training DEQ models (Bai et al., 2019). In the text that follows, for convenience, we always refer to a DEQ layer using its explicit function $g_{x}$ because, with the above restriction of $\mathrm{Fix}(g_{x})$ being a singleton, the implicit function $x\mapsto z^{\star}\in\mathrm{Fix}(g_{x})$ is well-defined.
One of the simplest numerical schemes for computing fixed points, which is the numerical scheme we consider in this study, is the standard fixed point iteration, which, using the notation in Definition 3.1, we can write as
@@ 3.2 Standard Interference and Positive Concave Mappings
    $(\forall x\in\mathbb{R}^{n}_{+})(\forall\lambda>1)\quad g(\lambda x)\ll\lambda g(x).$ (4)
@@
We emphasize that NC mappings are not SI mappings in general (e.g., $x\mapsto x$ for $x\in\mathbb{R}^{n}_{+}$ is an NC, but not an SI mapping), only PC mappings are guaranteed to be SI mappings without any further assumptions, see (Cavalcante et al., 2016, 2019).
@@ 4 PC Deep Equilibrium Layers
We recall that the activation functions from the above lists are applied elementwise to vector arguments, see Definition 3.1. The following lemma provides two simple sufficient conditions to construct positive concave DEQ layers.
@@
Consider a DEQ layer $g_{x}\colon\mathbb{R}^{n}_{+}\to\operatorname{int}(\mathbb{R}^{n}_{+})$ of the form in (1) in Definition 3.1 for an input $x$. Then:
  • (Assumption 1) $z\mapsto g_{x}(z):=\sigma(Wz+x)$ is a PC mapping $g_{x}\colon\mathbb{R}^{n}_{+}\to\operatorname{int}(\mathbb{R}^{n}_{+})$ if $z\in\mathbb{R}^{n}_{+}$, $W\in\mathbb{R}^{n\times n}_{+}$, $x\in\textup{int}(\mathbb{R}^{n}_{+})$, and $\sigma$ is constructed elementwise from any scalar activation function from List 1 in Remark 4.1;
  • (Assumption 2) $z\mapsto g_{x}(z):=\sigma(Wz+x)$ is a PC mapping $g_{x}\colon\mathbb{R}^{n}_{+}\to\operatorname{int}(\mathbb{R}^{n}_{+})$ if $z\in\mathbb{R}^{n}_{+}$, $W\in\mathbb{R}^{n\times n}_{+}$, $x\in\mathbb{R}^{n}_{+}$, and $\sigma$ is constructed elementwise from the scalar activation function from List 2 in Remark 4.1.
@@
On the other hand, if Assumption 2 is satisfied, we note that, in this case, $\sigma(0)\gg 0$, hence $(\forall z\in\mathbb{R}^{n}_{+})\ g_{x}(z)=(\sigma\circ f_{x})(z)\gg 0$, which implies $g_{x}$ is also a PC mapping in this case. \ $\blacksquare$
@@
Let $g_{x}\colon\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ be a DEQ layer of the form in (1) for a given input $x$. If $g_{x}$ (including the input $x$) satisfies Assumption 1 in Lemma 4.2, then $g_{x}$ is called a pcDEQ-1 layer. If $g_{x}$ satisfies Assumption 2 in Lemma 4.2, then $g_{x}$ is called a pcDEQ-2 layer. By pcDEQ we mean a layer $g_{x}$ that is either a pcDEQ-1 layer or a pcDEQ-2 layer.
@@
Figure 1: The visualization of the possible construction of pcDEQ layers. The symbols shown in the figures mean: +$\mathbb{R}^{n\times n}_{+}\ni W_{+}$ nonnegative weights, +$\mathbb{R}^{n}_{+}\ni z_{+}$ nonnegative vector of fixed point iteration, +$\mathbb{R}^{n}_{+}\ni x_{+}$ nonnegative input to the layer, $\textup{int}(\mathbb{R}^{n}_{+})\ni x_{++}$ positive input to the layer, $\sigma_{NC}$ nonnegative concave activation function (List 1 in Remark 4.1) and $\sigma_{PC}$ positive concave activation function (List 2 in Remark 4.1).
Lemma 4.2 asserts that pcDEQ layers are PC mappings for any allowed input $x$, which, in view of Proposition A.6, implies that they are also SI mappings. We can now use Definition A.2 to provide the form of the asymptotic mapping associated with a pcDEQ layer. The assertion of Proposition 4.4 follows from Proposition 11 in (Piotrowski et al., 2024), and, for completeness, we include the proof below.
@@
$g_{x,\infty}(z)=\lim_{p\to\infty}\sigma^{\prime}(Wpz+x)Wz.$ (7)
For activation functions composed from any scalar activation functions in Remark 4.1, we have $\lim_{p\to\infty}\sigma^{\prime}(Wpz+x)=0$, hence also
@@
Choose any input $x$ satisfying the conditions in Definition 4.3. From Lemma 4.2, $g_{x}$ is a PC mapping. Thus, from Proposition A.6, $g_{x}$ is also an SI mapping. From Corollary 4.5, the nonlinear spectral radius of $g_{x}$ is $\rho(g_{x,\infty})=0.$ If follows from Proposition A.4 that $g_{x}$ has a unique fixed point $z^{\star}.$ Then, from Proposition A.7 it follows that the fixed point iteration of $g_{x}$ in (2) converges geometrically to $z^{\star}$ with a factor $c\in[0,1)$ for any starting point $z_{1}\in\mathbb{R}^{n}_{+}.$blacksquare